# C3VDv2 - Colonoscopy 3D video dataset with enhanced realism

### v1

@@ Acknowledgments
Figure 2: Overview of the C3VDv2 data collection process. a. High-resolution 3D models of colons are digitally sculpted. b. Three-part molds are 3D printed. The high-resolution core insert (black) represents the colon lumen, while the outer shells (white) define the external contours of the phantom. c. The phantom casting process involves (1) painting silicone layer-by-layer onto the insert mold with the vasculature hand-painted on intermediate layers. The painted insert is then placed in the outer shells, and (2) silicone is poured through mold inlets to fill the space between the insert and outer shell, d. resulting in the final colon phantom. e. The imaging protocol involves rigidly attaching the colonoscope tip to a robotic arm for precise, repeatable, and pose-tracked maneuvers. Phantoms are secured inside their outer shells to maintain shape. f. 2D-3D registration is performed to align the 3D model with the 2D RGB video frames. The loss function is optimized based on the overlap of the corresponding edge maps. g. After alignment, a virtual camera is navigated along the camera trajectory to render pixel-wise ground truth frames.
Figure 3: Cross-section of Colon models: High-resolution 3D models were digitally sculpted referencing anatomical images. For positional reference, magnified views of lesions and Ileocecal valves are shown. The colors represent the distinct colon segments used for making phantoms. Quantitative details such as polyp sizes are provided in Table 2 and 3.
@@
Figure 6: Technical validation results: a) To verify the synchronization between the clean (v2) and debris-filled (v3) videos, the structural similarity index (SSIM) was computed across a window of frames with offsets. Maximum SSIM at offset 0 indicates that clean & debris-filled colon videos are well synchronized. b) The alignment of the RGB video frames and the depth maps was qualitatively and quantitatively evaluated. The quantitative score comprises a combination of dice score and chamfer distance, while the qualitative score was given by visual inspection of RGB video frames overlaid on rendered depth maps (1 - best, 3 - least). The qualitative scores agree with the quantitative scores. c) Histogram of challenging scenarios with respect to qualitative score. This indicates the presence of realistic scenarios, even in well-aligned videos with scores of 1 & 2.
Figure 7: Video snapshots of challenging scenarios: The C3VDv2 dataset features a wide range of challenging scenarios such as fast non-smooth motion, scope dipping to debris, lens cleaning, water on the lens, water jet, exploratory helical motion, loops, and en face to down the barrel transitions. Evaluating 3D reconstruction algorithms on C3VDv2 will provide a more representative assessment of their performance in real-world clinical settings. Green boxes indicate the same image.
@@
Camera type Res. FOV Tissue Type 3D GT Pose 3D GT Type Debris Blood Foam Water on Lens Deformation Frames Dataset
@@
Rendered SD Narrow Virtual CT Depth 23,421 SimCol3D [21, 22]
@@
Colonoscope HD Wide Phantom Sculpted Depth, Normals, Flow, 3D Model, Diffuse 67,886 C3VDv2 (Proposed)
Colonoscope HD Wide Phantom Sculpted Undeformed 3D Model 6,185 C3VDv2 (Proposed)
@@ B Ascending 164 - - - -
C Transverse 1 132 4 Adenoma IIa + IIs 10
D Transverse 2 160 - - - -
E Descending 149 5 Adenoma Is 3
F Sigmoid 1 150 6 Adenoma IIa 6.7
@@ H Rectum 167 - - - -
A Cecum 83 1 Ileocecal valve - 15.3
B Ascending 160 2 3 Adenoma Adenoma Ip IIa 5.7 11.4
C Transverse 1 148 4 Hyperplastic Is 5.1
D Transverse 2 100 5 Hyperplastic IIa 4.4
E Descending 170 6 Adenoma Is 3.9
F Sigmoid 94 7 8 Adenoma Hyperplastic Is IIa 3.3 11.5
G Rectum 105 - - - -

### v2

@@ Acknowledgments
Author contributions statement
M.V.G., T.L.B., and N.J.D. conceptualized and designed the study. M.V.G., L.S.G.F., and L.A. fabricated the phantoms, acquired and processed the data. T.L.B. implemented the registration and rendering pipeline. V.S.A. provided clinical oversight and performed simulated colonoscopies. M.V.G. performed the analysis and technical validation. M.V.G., L.A., L.S.G.F., and N.J.D. wrote the initial manuscript draft. All authors contributed to manuscript review and editing. N.J.D. supervised the project and acquired funding.
Competing interests
V.S.A. is the co-founder of Origin Endoscopy Inc., Sotelix Endoscopy Inc., Flow Therapeutics Inc. and Solv Endotherapy Inc., and a consultant for Dragonfly Endoscopy, Ionis Pharmacy, and Olympus Medical. V.S.A has educational grants from Boston Scientific and Medtronic, and research grants from Nestle and Pentax Medical. All other authors have no relevant competing interests.
@@
Figure 2: Video snapshots of challenging scenarios: The C3VDv2 dataset features a wide range of challenging scenarios such as fast non-smooth motion, scope dipping in debris, lens cleaning, water on the lens, water jet, exploratory motion, loops, and en face to down the barrel transitions. Evaluating 3D reconstruction algorithms on C3VDv2 will provide a more representative assessment of their performance in real-world clinical settings. Green boxes indicate the images from the same position at the end of the loop.
Figure 3: Overview of the C3VDv2 data collection process. (a) High-resolution 3D models of colons are digitally sculpted. (b) Three-part molds are 3D printed. The high-resolution core insert (black) represents the colon lumen, while the outer shells (white) define the external contours of the phantom. (c) The phantom casting process involves (1) painting silicone layer-by-layer onto the insert mold with the vasculature hand-painted on intermediate layers. The painted insert is then placed in the outer shells, and (2) silicone is poured through mold inlets to fill the space between the insert and outer shell, (d) resulting in the final colon phantom. (e) The imaging protocol involves rigidly attaching the colonoscope tip to a robotic arm for precise, repeatable, and pose-tracked maneuvers. Phantoms are secured inside their outer shells to maintain shape. (f) 2D/3D registration is performed to align the 3D model with the 2D RGB video frames. The loss function is optimized based on the overlap of the corresponding edge maps. (g) After alignment, a virtual camera is navigated along the camera trajectory to render pixel-wise ground truth frames.
Figure 4: Cross-section of Colon models: High-resolution 3D models were digitally sculpted referencing anatomical images. For positional reference, magnified views of lesions and Ileocecal valves are shown. The colors represent the distinct colon segments used for making phantoms. Quantitative details such as segment lengths and polyp sizes are provided in Table 2 and 3.
@@
Figure 7: Technical validation results: (a) To verify the synchronization between the clean (v2) and debris-filled (v3) videos, the structural similarity index (SSIM) was computed across a window of frames with offsets. Maximum SSIM at offset 0 indicates that clean & debris-filled colon videos are well synchronized. (b) The alignment of the RGB video frames and the depth maps was qualitatively and quantitatively evaluated. The quantitative score comprises a combination of dice score and chamfer distance, while the qualitative score was given by visual inspection of RGB video frames overlaid on rendered depth maps (1 - best, 3 - least). The qualitative scores agree with the quantitative scores. (c) Histogram of challenging scenarios with respect to qualitative score. This indicates the presence of realistic scenarios, even in well-aligned videos with scores of 1 & 2.
@@
Camera type Res. FOV Tissue Type 3D GT Source Pose Provided 3D GT Debris Blood Foam Water on Lens Deformation Frames Dataset
@@
Rendered SD Narrow Virtual CT Depth 37,800 SimCol3D [11, 12]
@@
Colonoscope HD Wide Phantom Sculpted Depth, Normals, Flow, 3D Model, Diffuse 67,886 C3VDv2 (Pixel-wise GT videos)
Colonoscope HD Wide Phantom Sculpted Undeformed 3D Model 6,185 C3VDv2 (Deformation videos)
Colonoscope HD Wide Phantom Sculpted 3D model 95,300 C3VDv2 (Sim. Screening videos)
Ground Truth (GT), Standard Definition (SD), High Definition (HD), Computed Tomography (CT), Optical Scan (OS).
@@ B Ascending 164 - - - -
C Transverse 2 132 4 Adenoma IIa + Is 10.2
D Transverse 1 160 - - - -
E Descending 149
F Sigmoid 1 150 5 Adenoma Is 6.7
@@ H Rectum 167 - - - -
A Cecum 83 1 Ileocecal valve - 15
B Ascending 160 2 Adenoma Ip 5.7
3 Adenoma Is 11.4
4 Hyperplastic IIa 6.2
C Transverse 1 148 5 Hyperplastic IIa 6.0
D Transverse 2 100 6 Adenoma Is 4.7
E Descending 170 - - - -
F Sigmoid 94 7 Adenoma Is 3.6
G Rectum 105 8 Hyperplastic Is 11.6